if the length of the common chord of the circle x2 + y2 = 4 and the parabola y2 = 4(x-h) is maximum, then find the value of h.

Grade:11

1 Answers

Deepak Kumar Shringi
5 years ago
Length will be maximum when LR of parabola will become diametere of circle that is line x=0
so in that case focus (0,0) here focus is x-h=1 and y=0 so x=1+h but we want zero so h=-1